OpenStudy (anonymous):

solve 27^x=9^(2x-7), show work and explain steps

OpenStudy (jdoe0001):

\(\large { 27^x=9^{2x-7}\qquad \begin{cases} 27\to 3\cdot 3\cdot 3\to 3^3 \\ \quad \\ 9\to 3\cdot 3\to 3^2 \end{cases}\qquad thus \\ \quad \\ 27^x=9^{2x-7}\implies (3^3)^x=(3^2)^{2x-7}\implies 3^{{\color{blue}{ 3x}}}=3^{{\color{blue}{ 2(2x-7)}}} }\)